a ship is travelling due east at 10 km/h. another ship heading 30degree east of north is always due north from the first ship.  the speed of second ship in km/h is

Grade:12th Pass

2 Answers

Sanjeev Malik IIT BHU
12 years ago

always solve by leting two components as v1 in north nad v2(in east....as net b/w north n east).

as original speed at 30deg east of north so, so perpendiculat to it zero velocity.

v1sin30= v2sin60                 .........................1

also by relative concept as relative to other ship dirn is north so zero vel along east west direction...it gives

v2=10

 

thus find the ans...

 

 

"like" the post below and click "yes" if it helps

DEVANSH
5 years ago

two components is v1 in north nad v2(in east....as net b/w north n east).

as original speed at 30deg east of north so, so perpendiculat to it zero velocity.

v1sin30= v2sin60                 .........................1

also by relative concept as relative to other ship dirn is north so zero vel along east west direction...it gives

v2=10

 

then x=10/sin(30º)=10/1/2=20 km/h